Please answer 2.4.c1 and 2.5.4
2.4.C1 You are driving and come to a fork in the road. Event A is the event that you turn left. Event B is the event that you turn right. A and B are mutually exclusive. True or False? ○ False
media%2Fb68%2Fb68174f5-cf16-44f5-bf6a-38
Answer #1

2.4 C1 True they are mutually exclusive as cant turn right and left at same time.
.....................
2.5.4

Evidence of Gas Leaks yes Evidence of electrical failurelyes 50 31 81 7 12 19 57 43 100

(a)
A failure involved a gar leak=50/100=0.5

(b)The there is evidence of electrical failure given that there was a gas leak
=50/57=0.87719
rounded =0.877

(c)That there is evidence of a gas leak given that there is evidecne of electrical failure
=50/81=0.6172

...................

if there is any doubt ask in comment give thumbs up if you like work.